A 3.556g sample of a pure aluminum oxide decomposes under heat to produce 1.674g of oxygen in addition to pure aluminum metal. What is the empirical formula of the aluminum oxide?
Answer #1

we have mass of each elements as:

Al: 3.556 - 1.674 = 1.882 g

O: 1.674 g

Divide by molar mass to get number of moles of each:

Al: 1.882/26.98 = 6.976*10^-2

O: 1.674/16.0 = 0.1046

Divide by smallest:

Al: 6.976*10^-2/6.976*10^-2 = 1

O: 0.1046/6.976*10^-2 = 1.5

Multiply by 2 to get simplest whole number ratio:

Al: 1*2 = 2

O: 1.5*2 = 3

So empirical formula is:Al2O3

Answer: Al2O3